# Lan Wang1,2∗        Yujia Chen2        Du Tran2       Vishnu Naresh Boddeti1       Wen-Sheng Chu2

### v2

@@ Appendix C Additional Qualitative Results
Appendix D More Implementation Details
@@ D.1 Token Extraction
Object token. For object tokens, we extract masks using from SAM2 [32] Automatic Mask Generator. For mask prediction, we sample $64\times 64$ points per image for dense and uniform coverage, with a batch size of 128 points to balance computational efficiency and memory usage. Predicted masks are filtered using a quality threshold of pred_iou_thresh=0.88, retaining only masks with high predicted IoU scores, and a stability score threshold of stability_score_thresh=0.92, ensuring the robustness of masks under varying binarization cutoffs. To calculate the stability score, the cutoff is shifted by stability_score_offset=0.99. Non-maximal suppression (NMS) is applied with an IoU threshold of box_nms_thresh=0.7 to remove redundant masks. We do not employ additional cropping layers (crop_n_layers=0). We extract features based on the mask’s bounding box and use the same encoder as the scene token for different datasets. We set $N_{\text{key}}=128$ for MovieChat and $N_{\text{key}}=64$ for Ego4D-NLP and LVBench datasets.
In Attention Learning stage, we sample in total 256 tokens for MovieChat, 200 / 450 tokens for Ego4D-NLP and 16 tokens for LVBench.

### v3

@@ Appendix C Additional Qualitative Results
Appendix D Additional Implementation Details
@@ D.1 Token Extraction
Object token. For object tokens, we extract masks using from SAM2 [32] Automatic Mask Generator. For mask prediction, we sample $64\times 64$ points per image for dense and uniform coverage, with a batch size of 128 points to balance computational efficiency and memory usage. Predicted masks are filtered using a quality threshold of pred_iou_thresh=0.88, retaining only masks with high predicted IoU scores, and a stability score threshold of stability_score_thresh=0.92, ensuring the robustness of masks under varying binarization cutoffs. To calculate the stability score, the cutoff is shifted by stability_score_offset=0.99. Non-maximal suppression (NMS) is applied with an IoU threshold of box_nms_thresh=0.7 to remove redundant masks. We do not employ additional cropping layers (crop_n_layers=0). We extract features based on the mask’s bounding box, expand it by 2x to include additional contextual information, and use the same encoder as the scene token for different datasets. We set $N_{\text{key}}=128$ for MovieChat and $N_{\text{key}}=64$ for Ego4D-NLP and LVBench datasets.
@@
In Attention Learning stage, we sample in total 256 tokens for MovieChat, 200 / 450 tokens for Ego4D-NLP and 16 tokens for LVBench. Note that since the task of Ego4D-NLP is temporal grounding, we only used action tokens and scene tokens to ensure temporal consistency.